OpenStudy (anonymous):

The length of a rectangle is equal to three times its width. If the perimeter is equal to 240 feet, what is the length of the rectangle?

OpenStudy (misssunshinexxoxo):

w=240/8, =30 0

OpenStudy (misssunshinexxoxo):

Then we multiply l=3w=3*30=90

OpenStudy (anonymous):

beat me to it :P

OpenStudy (misssunshinexxoxo):

@toturnt57 do you understand?